OpenStudy (nuccioreggie):

A frog begins at negative twenty-one on a number line. The frog hops positive sixteen times and then hops negative four times. At what value does the frog end? negative forty-one negative nine positive nine negative one

OpenStudy (nurali):

-25 + 16 = -9

OpenStudy (anonymous):

I have arrived at the same conclusion as Nurali, but through a different way. -21 + 16 = -5 -4 =-9.